OpenStudy (anonymous):

In a coordinate system, points A(-2, 3) and B(6, -1) have a midpoint of X. What is the midpoint between A and X?

OpenStudy (anonymous):

coordinates are written as (x,y). first find the mid point in x axis- (-2 and 6) ((6)+(-2))/2=2.... then they y coordinates ((3)+(-1))/2=1 so the answer is (2,1)

OpenStudy (anonymous):

thanks alot

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ A&(-2\quad ,&3)\quad B&(6\quad ,&-1) \end{array} \\\quad \\ \text{middle point of 2 points }\\ X=\left(\cfrac{x_2 + x_1}{2}\quad ,\quad \cfrac{y_2 + y_1}{2} \right) \)

OpenStudy (anonymous):

so its the formula for midpoint?...thnkss...its getting easier :)

OpenStudy (anonymous):

A rectangular display case houses a square pyramid. Both have the same square base measuring x inches on a side. The top of the case is 2 inches above the tip of the pyramid which has a height of h inches. Which expression represents the volume of the empty space inside the case?